OpenStudy (anonymous):

Find the x and y intercept : X/3-y/7=5

OpenStudy (anonymous):

put x=0 to get y-intercept put y=0 to get x-intercept

OpenStudy (anonymous):

?

OpenStudy (ghazi):

\[\frac{ x }{ a }+\frac{ y }{ b }=1\] a and b are intercepts ..so divide your equation by 5

OpenStudy (ghazi):

\[\frac{ x }{ 5*3 }+\frac{ y }{ 7*5 }=1\] hope x and y are clear to you

OpenStudy (ghazi):

i mean x and y intercepts

OpenStudy (ghazi):

@NatalieLove is it clear?

hartnn (hartnn):

@ghazi *minus*

OpenStudy (ghazi):

@hartnn sorry it is what i was gonna edit lol

OpenStudy (anonymous):

equation line where intercept with x-axis is known as x-intercept and similarly y-intercept

OpenStudy (anonymous):

What is. The solution

OpenStudy (ghazi):

it is in front of you

OpenStudy (anonymous):

x=15,y=35

OpenStudy (ghazi):

\[\frac{ x }{ 5*3 }+\frac{ y }{ -7*5 }=1...a=5*3,.... b=-7*5\] is that clear